# Spatial vertical regression for spatial panel data: Evaluating the effect of the Florentine tramway’s first line on commercial vitality

### v2

@@ 6.2 Alternative approaches
To implement the original SC method (Abadie et al., 2010), we adopt Euclidean distance for the estimation of the weights. Furthermore, we implement three vertical regression methods (Doudchenko and Imbens, 2016) using the simple Ordinary Least Squares estimator (OLS), the ridge regression estimator (SR; which stands for “separable ridge” since we implement the ridge regression estimator for each treated unit separately), and a Bayesian Vertical Regression (BVR) approach. Finally, we implement a Bayesian synthetic control method (BSC, implemented in Stan; e.g. Kim et al., 2020; Martinez and Vives-i Bastida, 2022) using a Bayesian simplex regression approach. See Supplement 1 for the choice of prior distributions. For the Bayesian methods, we consider three chains of $10,000$ iterations each and a burn-in of 5,000. We assess the convergence using the potential scale-reduction statistic (Gelman and Rubin, 1992).
@@ 6.3 Evaluation metrics
We opt for these four methods as there are the ones with a clear quantification of the uncertainty.
@@ 6.4 Simulation results
SVR 0.00 0.00 -0.01
Table 1: Average post-treatment bias of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Spatial Vertical Regression (SVR) in scenarios with spatially mid-correlated weights ($\rho^{2}_{s}=0.4^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.34 0.41 0.68
Table 2: Average post-treatment MSE of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Vertical Regression (SVR) in scenarios with spatially mid-correlated weights ($\rho^{2}_{s}=0.4^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Table 3: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Bayesian Spatial Vertical Regression (SVR) in scenarios with spatially mid-correlated weights ($\rho^{2}_{s}=0.4^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Tables 1 and 2 show the average bias and MSE for SVR and the five alternative methods, respectively, and Table 3 shows ACP for OLS vertical regression, BVR, BSC and our SVR method. Unsurprisingly, all the methods perform well in terms of bias under all scenarios. At the same time, SVR has the smallest MSE in almost all scenarios, illustrating that it leads to more efficient imputations of the missing control potential outcome in the presence of spatial correlation in the controls’ coefficients. The only exceptions are a subset of scenarios with short pre-intervention period ($T_{0}=10$) and spatial residual errors. Here, SC and SR perform slightly better, possibly because regularization helps deal with precision problems (Doudchenko and Imbens, 2016). However, SVR leads to straightforward and appropriate uncertainty quantification, as we can see in Table 3. The ACP of the posterior 95% credible intervals using SVR is very close to the nominal level of 95% in all scenarios. In contrast, BVR tends to provide credible intervals that are relatively wide, over-covering the true values, and OLS and BSC have coverage of the 95% intervals that is substantially lower than their nominal level.
@@
Additional simulation results are included in Supplement 4, including results for the other levels of spatial correlation across the control units’ coefficients, $\rho^{2}_{s}$. We summarize the results here. Conclusions remain overall similar, with SVR performance being stronger against alternatives for larger values of spatial correlation, as expected. At the same time, its performance is comparable to alternative methods even in the absence of spatial correlation ($\rho^{2}_{s}=0.001^{2}$), especially for larger values of $T_{0}$. Lastly, Figure 3 reports the MSE for SVR estimates when $T_{0}=20$ varying the spatial correlation, $\rho^{2}_{s}$, the length of the post-treatment period and the error correlation structure. Evidently, the MSE of SVR decreases with higher level of spatial correlation under all the scenarios considered, illustrating that SVR’s performance improves in the presence of spatial correlation across the coefficients. On the other hand, the alternative methods do not exploit the spatial structure to reduce MSE (see Figures S.17, S.17, S.18 in Supplement 4).
@@
Overall, our simulation results show that SVR performs well under different scenarios. In scenarios where the presence of spatial correlation implies that weights defining the control potential outcomes vary smoothly across spatially correlated treated units, SVR generally outperforms the competitors, improving the precision of inferences in terms of both MSE and coverage probability of corresponding intervals. At the same time, it performs similarly to the competitors even in scenarios where there is no spatial correlation across treated units or there is only residual spatial correlation. Lastly, SVR leads to good inferential results throughout all cases considered, a challenging task in the synthetic control literature.
@@ 7 Estimating the effect of the Florentine tramway construction
Reported results are on the original scale. In Supplement 3 we include the time series and correlation plot for the treated and control outcomes after normalization and de-trending.
@@
Figure 4: Estimated $\bm{\beta}_{c}$ coefficients for three randomly chosen control units from three different estimation methods. See Figure S.7 in Supplement 3 for the figures corresponding to all control units.
Figure 5: Posterior medians (lines) and 95% posterior credible intervals (shaded areas) of the treatment effects $\Delta_{it}$ for areas within $d-$seconds of walking from a tramway stop, $d=120,240,360,480,600$. First vertical line: tramway worksite starts (2006) - Second vertical line: tramway operational (2010).
Figure 5 and Table 4 show the results using SVR. Figure 5 shows the posterior medians and 95% posterior credible intervals for the effects $\Delta_{it}$ defined in Equation 3, where intervals are constructed using the 2.5 and 97.5 quantiles of the posterior distribution. Table 4 reports the same quantities for the intertemporal average causal effects $\Delta_{i}$ defined in Equation 4, where the average is defined separately over the construction phase (2006–2009) and the operational phase (2010–2016).
@@
Estimate -0.260 -2.692 -7.652 -0.702 -3.692
CI (-6.434, 6.374) (-12.473, 7.918) (-23.236, 9.646) (-15.111, 15.006) ( -27.991, 23.162)
RMSPE 0.466 1.054 1.096 1.394 1.038
First of all, Figure 5 suggests that SVR provides a good fit for the data, since the estimates and 95% intervals in the pre-treatment period well-overlap with zero. During the construction phase, our results show positive effects for the area nearest to the tramway stops ($\leq$2 minute-walk) which attenuate at larger distances, though the credible intervals always cover zero. However, during the operational phase of the tramway line, we estimate a sharp decline in the number of shops immediately following the opening of the line in 2010 and 2011, followed by a rebound and growth in the subsequent years. This pattern can be explained by the economic environment in Italy during those years. The 2008 economic crisis and the problematic state of Italian sovereign debt slowed the opening of new shops and commercial hubs everywhere in Florence. In the tramway area specifically, the effect might have been more pronounced than elsewhere, given its ongoing renewal of commercial offerings at precisely the time the crisis hit. As a result, the overall impact of the tramway’s construction was delayed, although it becomes noticeable again starting in 2013. However, the credible intervals almost always cover zero, so we cannot draw firm conclusions.
During the operational phase of the tramway, qualitative conclusions across areas located at at different distances from a tram stop are similar. During the construction period, there is an increase in the number of stores in the area closest to the tramway (within 120 seconds of the stops) which is not evident at farther distances. This illustrates that businesses might have invested in the tramway area before its operational phase. These results are consistent with Grossi et al. (2025) where they found small effects for stores that sell durable goods and were located in the close proximity of the tramway line, before the tramway became operational.
@@ 8 Concluding Remarks
Supplement
1 The choice of hyperparameters of the prior distributions
1.1 Prior distributions in SVR
@@
  • For the parameters $b_{c}$ which represent the overall importance of a control unit across radii, we have specified $b_{c}\sim N(\mu_{b},\sigma^{2}_{b})$ with $\mu_{b}=0,\sigma^{2}_{b}=1$. Therefore, a priori the overall importance of a control unit is between $-3$ and 3 with probability 99.7%. This is reasonable considering that all data are standardized. The predictor (outcomes of the control unit) and the outcome (outcomes of the treated unit) in the vertical regression have standard deviation 1, therefore it is expected that the corresponding coefficients are not too large. A similar argument holds for the intercepts.
  • The parameters $\sigma^{2}_{\beta}$ is the variance of the values in $\bm{\beta}_{c}$ around their mean $b_{c}$. We specify an Inverse-Gamma prior on $\sigma^{2}_{\beta}$ and set hyperparameters for the shape $\alpha_{\beta}=2.7$ and the scale $\eta_{\beta}=0.3$. This choice implies that $P(\sigma^{2}_{\beta}<1)\approx 0.99$ and $\textrm{E}(3\sigma_{\beta})\approx 1.17$. Therefore, since the entries of $\bm{\beta}_{c}$ are normally distributed around $b_{c}$ with variance $\sigma^{2}_{\beta}$, they fall within $-3\sigma_{\beta}$ and $3\sigma_{\beta}$ around $b_{c}$ with probability 99.7%, which is specified to be most of the time between $-3$ and 3, and on average from $-1.17$ to 1.17.
  • The spatial range of the coefficients $\bm{\beta}_{c}$ is dictated by $\rho^{2}_{\beta}$ on which we impose a Gamma prior with shape parameter $\alpha_{\rho_{\beta}}=0.5$ and rate parameter $\eta_{\rho_{\beta}}=2$. This prior specifies that a priori $P(\rho^{2}_{\beta}<0.1)\approx 0.47,P(\rho^{2}_{\beta}<0.3)\approx 0.73,P(\rho^{2}_{\beta}<0.5)\approx 0.83$, and $P(\rho^{2}_{\beta}<1)~{}95\%$. Considering that all distances are standardized to cover $[0,1]$, this prior assigns sufficient weight to all reasonable values of $\rho^{2}_{\beta}$.
In Figure S.2, we include the curves for 30 independent draws from the induced prior distribution on $\bm{\beta}_{c}$. We see that the values cover a range from at least $-3$ to 3, and these draws include curves with strong or essentially absent spatial correlation.
@@
  • For the overall variance of the residuals we have specified the following shape and scale parameters in the Inverse-Gamma prior: $a_{e}=2$, and $\eta_{e}=0.35$. This distribution specifies that, a priori, $\sigma^{2}_{e}$ is less than 0.1, 0.3, 0.5, and 1 with probability that is approximately equal to 0.13, 0.67, 0.85, and 0.96, respectively. Since the outcome is standardized to have (unconditional) variance equal to 1, it is preferable that the prior distribution for the residual variance is concentrated on values below 1. Furthermore, this prior allows for a wide range of values for the residual variance, implying that the control units’ outcomes can explain little or a lot of the variability in a treated unit’s outcomes.
@@
For illustration, we consider 20 treated areas with values $d_{i}$ that are equally spread from 0 to 1, and we draw the matrix $\Sigma_{e}$ from its induced prior distribution 1,000 times. Figure S.2 shows the histogram of the induced correlation of the treated area with $d_{i}=0$ with the areas with the remaining values of $d_{i}$, across the 1,000 draws. We see that the induced prior distribution allows for essentially non-existent spatial correlation even at the smallest distances, while also allowing for a spatial correlation that is strong at moderate distances and declines at larger distances. Therefore, the induced prior covers a wide spectrum of spatial structures.
1.2 Prior distributions in BVR and BSC
For BVR, we use standard Normal prior on the regression coefficients and an Inverse Gamma distribution with shape and scale parameters equal to 4 and 2 respectively for the variance of the residuals. Similarly to the justification of SVR, this choice of hyperparameters is made since it satisfies that the probability that the variance is below 1 is, a priori, approximately 85%. For BSC the prior specification is the same, while we constrain the regression coefficients to lie in the simplex.
2 Illustration of simulation specifications

### v3

@@ 6.2 Alternative approaches
To implement the original SC method (Abadie et al., 2010), we adopt Euclidean distance for the estimation of the weights. Furthermore, we implement three vertical regression methods (Doudchenko and Imbens, 2016) using the simple Ordinary Least Squares estimator (OLS), the ridge regression estimator (SR; which stands for “separate ridge” since we implement the ridge regression estimator for each treated unit separately), and a Bayesian Vertical Regression (BVR) approach. Finally, we implement a Bayesian synthetic control method (BSC, implemented in Stan; e.g. Kim et al., 2020; Martinez and Vives-i Bastida, 2022) using a Bayesian simplex regression approach. See Supplement A for the choice of prior distributions. For the Bayesian methods, we consider three chains of $10,000$ iterations each and a burn-in of 5,000. We assess the convergence using the potential scale-reduction statistic (Gelman and Rubin, 1992).
@@ 6.3 Evaluation metrics
We opt for these four methods as they are the ones with a clear quantification of the uncertainty.
@@ 6.4 Simulation results
SVR 0.00 0.00 0.00
Table 1: Average post-treatment bias of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially mid-correlated weights ($\rho^{2}_{s}=0.4^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods. Bias and outcome mean value are expressed in the same unit of measure.
@@
SVR 0.33 0.40 0.66
Table 2: Average post-treatment MSE of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially mid-correlated weights ($\rho^{2}_{s}=0.4^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure: IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Table 3: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially mid-correlated weights ($\rho^{2}_{s}=0.4^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure: IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Tables 1 and 2 show the average bias and MSE for SVR and the five alternative methods, respectively, and Table 3 shows ACP for OLS vertical regression, BVR, BSC and our SVR method. Unsurprisingly, all the methods perform well in terms of bias under all scenarios. At the same time, SVR has the smallest MSE in all scenarios with mid spatial correlation ($\rho^{2}=0.4^{2}$), illustrating that it leads to more efficient imputations of the missing control potential outcome in the presence of spatial correlation in the controls’ coefficients. Notably, this result is confirmed for a short pre-intervention period ($T_{0}=10$), which is very common in applied studies.
At the same time, SVR leads to straightforward and appropriate uncertainty quantification, as we can see in Table 3. The ACP of the posterior 95% credible intervals using SVR is very close to the nominal level of 95% in all scenarios. BVR yields similar ACP results, but a higher MSE across all scenarios. In contrast, OLS and BSC have coverage of the 95% intervals that is substantially lower than their nominal level.
@@
Additional simulation results are included in Supplement D, including results for the other levels of spatial correlation across the control units’ coefficients, $\rho^{2}_{s}$. We summarize the results here. Conclusions remain overall similar, with SVR performance being stronger against alternatives for larger values of spatial correlation, as expected. SVR has lower MSE comparing with alternative methods, excluding the scenario with short pre-treatment, low spatial correlation across outcomes, and high error, where SR has marginally lower MSE. At the same time, its performance is comparable to alternative methods even in the absence of spatial correlation ($\rho^{2}_{s}=0.001^{2}$), especially for larger values of $T_{0}$. Lastly, Figure 3 reports the MSE for SVR estimates when $T_{0}=20$ varying the spatial correlation, $\rho^{2}_{s}$, the length of the post-treatment period and the error correlation structure. Evidently, the MSE of SVR decreases with higher level of spatial correlation under all the scenarios considered, illustrating that SVR’s performance improves in the presence of spatial correlation across the coefficients. On the other hand, the alternative methods do not exploit the spatial structure to reduce MSE (see Figures S.21, S.21, S.22 in Supplement D).
@@
Overall, our simulation results show that SVR performs well under different scenarios. In scenarios where the presence of spatial correlation implies that weights defining the control potential outcomes vary smoothly across spatially correlated treated units, SVR outperforms the competitors, improving the precision of inferences in terms of MSE and coverage probability of corresponding intervals. At the same time, it performs similarly to the competitors even in scenarios where there is no spatial correlation in the regression coefficients across treated units. Lastly, SVR leads to good inferential results throughout all cases considered, a challenging task in the synthetic control literature.
@@ 7 Estimating the effect of the Florentine tramway construction
Reported results are on the original scale. In Supplement C we include the time series and correlation plot for the treated and control outcomes after normalization and de-trending.
@@
Figure 4: Estimated $\bm{\beta}_{c}$ coefficients for three randomly chosen control units from three different estimation methods. See Figure S.7 in Supplement C for the figures corresponding to all control units.
Figure 5: Posterior medians (lines) and 95% posterior credible intervals (shaded areas) of the treatment effects $\Delta_{it}$ for areas within $d$-seconds of walking from a tramway stop, $d=120,240,360,480,600$. The two vertical lines represent the beginning of the tramways construction and operational phases in 2006 and 2010, respectively.
Figure 5 and Table 4 show the results using SVR. Figure 5 shows the posterior medians and 95% posterior credible intervals for the effects $\Delta_{it}$ defined in Equation 3, where intervals are constructed using the 2.5 and 97.5 quantiles of the posterior distribution. Table 4 reports the same quantities for the intertemporal average causal effects $\Delta_{i}$ defined in Equation 4, where the average is defined separately over the construction phase (2006–2009) and the operational phase (2010–2016). Results based on the alternative methods evaluated in Section 6 are reported in Figure S.9 in Supplement C.
@@
Estimate 1.803 0.359 -2.510 3.228 3.650
CI (-2.938, 6.485) (-7.076, 7.839) (-14.645, 10.008) (-8.065, 15.378) (-14.930, 23.361)
RMSPE 0.834 1.584 1.291 2.094 2.769
First of all, Figure 5 suggests that SVR provides a good fit for the data, since the estimates and 95% intervals in the pre-treatment period well-overlap with zero. Moreover, the overall RMSPE is ranging between 0.6% and 2.2% of the pre-treatment average outcome, suggesting a reasonable fit. During the construction phase, our results show positive effects for the area nearest to the tramway stops ($\leq$2 minute-walk) which attenuate at larger distances, though the credible intervals always cover zero. This illustrates that businesses might have invested in the tramway area during the construction phase, before the tramway became operational. However, during the operational phase of the tramway line, we estimate a sharp decline in the number of shops immediately following the opening of the line in 2010 and 2011, followed by a rebound and growth in the subsequent years. This pattern can be explained by the economic environment in Italy during those years. The 2008 economic crisis and the problematic state of Italian sovereign debt slowed the opening of new shops and commercial hubs everywhere in Florence. In the tramway area specifically, the effect might have been more pronounced than elsewhere, given its ongoing renewal of commercial offerings at precisely the time the crisis hit. As a result, the overall impact of the tramway’s construction was delayed, although it becomes noticeable again starting in 2013. However, the credible intervals almost always cover zero, so we cannot draw firm conclusions.
We further study how the impact of the tramway on commercial vitality varies across types of stores, particularly regarding the purchase frequency of the products offered. Following Grossi et al. (2025), we distinguish between stores selling durable and non-durable goods. We find that stores selling durable goods exhibit limited sensitivity to the tramway, with positive effects only within a large area from the tram stops in the later post-treatment period. In contrast, non-durable goods stores experience a temporary contraction during construction, followed by a clear post-construction increase in activity, particularly within intermediate walking distance from the tram stops. We present the results in Supplement C.3.
@@ 8 Concluding Remarks
Appendix A The choice of hyperparameters of the prior distributions
A.1 Prior distributions in SVR
@@
  • For the parameters $b_{c}$ which represent the overall importance of a control unit across radii, we have specified $b_{c}\sim\hbox{Laplace}(\lambda_{b})$ with $\lambda_{b}=0.1$. Therefore, a priori the overall importance of a control unit is between 1-1 and 1 with probability 99.97%. This is reasonable considering that all data are standardized. Since outcomes are standardized to a unit standard deviation, the coefficients of the control units in the vertical regression model are expected to be limited in magnitude.
  • The parameter $\sigma^{2}_{\beta}$ represents the variance of the coefficients in $\boldsymbol{\beta}_{c}$ around their mean $b_{c}$. We specify a Truncated Normal prior in zero to $\sigma^{2}_{\beta}$ with hyperparameters $(a_{\beta},\eta_{\beta})=(0,\,0.35)$. This choice corresponds to a commonly used weakly informative prior, [Gelman, 2006] which exerts minimal influence on the scale of $\boldsymbol{\beta}_{c}$ while still allowing the model to regularise extreme values of the variance. Because the prior places substantial mass near zero and exhibits very heavy tails, it encourages the coefficients to remain concentrated around their centre unless the data provide strong evidence for a larger variance. Similarly, the error variance $\sigma^{2}_{e}$ is assigned an Truncated Normal prior in zero prior with hyperparameters $(a_{e},\eta_{e})=(0,\,0.5)$, which again yields a weakly informative and scale–flexible prior.
  • The spatial range of the coefficients $\bm{\beta}_{c}$ is dictated by $\rho^{2}_{\beta}$ on which we impose a Gamma prior with shape parameter $\alpha_{\rho_{\beta}}=0.5$ and rate parameter $\eta_{\rho_{\beta}}=2$. This prior specifies that a priori $P(\rho^{2}_{\beta}<0.1)\approx 0.47,P(\rho^{2}_{\beta}<0.3)\approx 0.73,P(\rho^{2}_{\beta}<0.5)\approx 0.83$, and $P(\rho^{2}_{\beta}<1)\approx 95\%$. Considering that all distances are standardized to cover $[0,1]$, this prior assigns sufficient weight to all reasonable values of $\rho^{2}_{\beta}$.
In Figure S.2, we include the curves for 30 independent draws from the induced prior distribution on $\bm{\beta}_{c}$. We see that the values cover a range from at least 1-1 to 1, and these draws include curves with strong or essentially absent spatial correlation.
@@
  • For the overall variance of the residuals we have specified the following shape and scale parameters in the Truncated Normal prior: $a_{e}=0$, and $\eta_{e}=0.5$. This distribution specifies that, a priori, $\sigma^{2}_{e}$ is less than 0.1, 0.3, 0.5, and 1 with probability that is approximately equal to 0.47, 0.73, 0.84, and 0.95, respectively. Since the outcome is standardized to have (unconditional) variance equal to 1, it is preferable that the prior distribution for the residual variance is concentrated on values below 1. Furthermore, this prior allows for a wide range of values for the residual variance, implying that the control units’ outcomes can explain little or a lot of the variability in a treated unit’s outcomes.
@@
For illustration, we consider 20 treated areas with values $d_{i}$ that are equally spread from 0 to 1, and we draw the matrix $\Sigma_{e}$ from its induced prior distribution 1,000 times. Figure S.2 shows the histogram of the induced correlation of the treated area with $d_{i}=0$ with the areas with the remaining values of $d_{i}$, across the 1,000 draws. We see that the induced prior distribution allows for spatial correlation at the smallest distances, while also allowing for a spatial correlation at moderate distances and declines at larger distances. Therefore, the induced prior covers a wide spectrum of spatial structures.
A.2 Prior distributions in BVR and BSC
For BVR, we specify a standard normal prior for the regression coefficients and a truncated normal distribution at zero with parameters 0 and $0.5$ for the residual variance. Similarly to the justification of SVR, this choice of hyperparameters is made since it satisfies that the probability that the variance is below 1 is, a priori, approximately 85%. For BSC the prior specification is the same, while we constrain the regression coefficients to lie in the simplex.
Appendix B Illustration of simulation specifications